OpenStudy (paystub-2):

I need to use graphing,substitution, or elimination: y=6x+1 y=7x

OpenStudy (anonymous):

using substitution sub y=7x into y=6x+1 to get 7x=6x+1 then solve for x=__?

OpenStudy (anonymous):

for graphing plot them using different x values closer to the value of what you've solved i.e. x=-2,-1,0.1,2 for y=7x y=7(-2)=__? y=7(-1)=___? y=7(0)=_____? y=7(1)=___? y=7(2)=___? etc.... and for y=6(-2)+1=___? y=6(-1)+1 =____? y=6(0)+1 =____? y=6(1)+1 =___? y=6(2)+1 =___? y=6(3)+1 =____? etc